OpenStudy (exploradora1011):

Can someone explain how to do this problem? Find the equation of the line specified. The slope is 6, and it passes through ( -4, 4).

Mehek (mehek14):

general equation y = mx + b m = slope b = y-intercept you are give slope (m) = 6 y = 6x + b plug in the ordered pair 4 = 6 * 4 + b Solve for b

OpenStudy (exploradora1011):

Thank you so much!

Mehek (mehek14):

You're welcome ^_^